# CRAG - Comprehensive RAG Benchmark

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has recently emerged as a promising solution to alleviate Large Language Model (LLM)’s deficiency in lack of knowledge. Existing RAG datasets, however, do not adequately represent the diverse and dynamic nature of real-world Question Answering (QA) tasks. To bridge this gap, we introduce the Comprehensive RAG Benchmark (CRAG), a factual question answering benchmark of 4,409 question-answer pairs and mock APIs to simulate web and Knowledge Graph (KG) search. CRAG is designed to encapsulate a diverse array of questions across five domains and eight question categories, reflecting varied entity popularity from popular to long-tail, and temporal dynamisms ranging from years to seconds. Our evaluation on this benchmark highlights the gap to fully trustworthy QA. Whereas most advanced LLMs achieve $\leq 34\%$ accuracy on CRAG, adding RAG in a straightforward manner improves the accuracy only to 44%. State-of-the-art industry RAG solutions only answer $63\%$ questions without any hallucination. CRAG also reveals much lower accuracy in answering questions regarding facts with higher dynamism, lower popularity, or higher complexity, suggesting future research directions. The CRAG benchmark laid the groundwork for a KDD Cup 2024 challenge, attracting thousands of participants and submissions within the first 50 days of the competition. We commit to maintaining CRAG to serve research communities in advancing RAG solutions and general QA solutions.
@@ 1 Introduction
Large Language Models (LLMs) have transformed the landscape of Natural Language Processing (NLP) tasks, especially in Question Answering (QA) [27, 15, 14]. Despite the advancements, the issue of hallucination persists as a significant challenge; LLMs may generate answers that lack factual accuracy or grounding [19, 9]. Studies have shown that GPT-4’s accuracy in answering questions referring to slow-changing or fast-changing facts is below 15% [25]; even for stable (never-changing) facts, GPT-4’s accuracy in answering questions referring to torso-to-tail (less popular) entities is below 35% [21]. Overcoming hallucinations thus becomes a priority in building reliable QA systems [9, 8].
@@
Retrieval-Augmented Generation (RAG) [7, 13, 5] has recently emerged as a promising solution to alleviate LLM’s deficiency in lack of knowledge and attracted a lot of attention from both academia research and industry. Given a question, a RAG system searches external sources to retrieve relevant information and then provides grounded answers [7, 13, 6] (see Figure 1 for an illustration). Despite its potential, RAG still faces many challenges, such as selecting the most relevant information, reducing question answering latency, and synthesizing information to answer complex questions.
A comprehensive benchmark is currently missing to advance continued research efforts in this field. Traditional QA benchmarks such as Natural Questions [12], TriviaQA [10], and MS MARCO [3] have advanced QA in the past decade but do not adequately represent the diverse and dynamic challenges RAG is facing. New benchmarks that focus on LLM or RAG, such as FreshQA [25] and RGB [5], usually target certain capabilities of LLMs and contain only a few hundred queries. The goal of our work is to build a comprehensive benchmark to propel the area forward.
What is a good benchmark for QA over LLMs? We consider five critical features.
@@
Our second contribution is the evaluation mechanism to allow for reliable comparisons. We designed 3 tasks to test different components in RAG solutions: web retrieval, structured querying, and summarization (Section 2). Instead of computing the percentage of correctly answered questions, our score system distinguishes hallucinated answers and missing answers, and gives the former a higher penalty as it can be more harmful to ruin user trust. We also design an effective automatic evaluation mechanism to allow for fast evaluations and iterations (Section 4).
Our third contribution is a comprehensive evaluation of straightforward RAG solutions and industry state-of-the-art solutions on RAG (Section 5). Whereas most advanced LLMs achieve $\leq 34\%$ accuracy on CRAG, adding RAG in a straightforward manner improves the accuracy only to 44%. State-of-the-art industry RAG solutions answer only $63\%$ questions without any hallucination, still having much lower accuracy in answering questions regarding facts with higher dynamism, lower popularity, or higher complexity. These evaluations serve two roles: first, they demonstrate that CRAG has appropriate level of difficulty, and allows insights drawn from different dimensions of diversities the benchmark has incorporated; second, they highlight the gaps and research directions to a fully trustworthy QA system.
@@
Comparison with existing benchmarks. Table 1 compares CRAG with existing benchmarks for factual question answering, which illustrates several advantages of the CRAG benchmark: comprehensive coverage, realistic testing with mock APIs, dynamic question handling, diverse fact popularity, and extensive beyond Wikipedia.
These features make CRAG a robust and versatile benchmark for testing RAG systems and broadly QA systems, providing a shared testbed in evaluating how these systems handle real-world, dynamic, and diverse information retrieval and synthesis challenges for reliable LLM-based question answering.
@@
Benchmark Web retrieval KG search Mock API Dynamic question Torso and tail facts Beyond Wikipedia
QALD-10 [24]
MS MARCO [3] not explicitly not explicitly
Natural Questions [12] not explicitly not explicitly
RGB [5]
FreshLLM [25]
CRAG
@@ 2 Problem Description
The three tasks, each adding upon the previous one, allow testing different capabilities of the end-to-end RAG systems.
@@ 3 Dataset Description
Post-processing heavy Questions that need reasoning or processing of the retrieved information to obtain the answer (e.g., “how many days did Thurgood Marshall serve as a Supreme Court justice?”).
@@
CRAG contains two parts of data: the QA pairs and the contents for retrieval. We now describe each part of the data. Data generation details can be found in Appendix A.1.1-A.1.6.
@@ 3.1 Question answering pairs
Using the above methods, we collected 2,425 Web Questions and 1,984 KG Questions, with $661$, $658$, and $665$ KG Questions containing head, torso, and tail entities respectively. Tables 3 and 4 summarize the distribution of the questions across different dimensions. The size of each dimension slice (e.g., fast-changing facts) allows us to get metrics with $<5\%$ margin-of-error (with 95% confidence level) for most of the cases. The dynamism distribution roughly reflects the nature of the domain (e.g., much more real-time questions for Finance than for other domains). See Appendix A.1.2 for definition of the dynamism categories.
@@ 3.2 Contents for retrieval
Web search results. For each question, we used the question text as the search query and stored up to $50$ HTML pages from the Brave search API [4]. See Table 8 in Appendix A.1.4 for an example. We estimated the web search recall ($50$ web pages) with a heuristic-based method: first check whether the ground truth answer URL was found among the 50 pages; if not, search whether the fact in the ground truths is contained in the page snippet or content. The estimated recall is $84\%$ for Web Questions and $63\%$ for KG Questions (Table 9 in Appendix A.1.5), consistent with our intuition that torso and tail entities in the KG questions may not be included in the 50 returned pages.
@@
In total, the resulting data contains 220K webpages, a KG of 2.6M entities, and 38 Mock APIs.
@@ 4.1 Metrics
We then use a scoring method Scoreh with score $1$, $0.5$, 00, and $-1$ for each perfect, acceptable, missing, and incorrect answer, respectively, where we penalize hallucinated answers and prefer missing answers to incorrect ones. For a given RAG system, we compute the average score from all examples in the evaluation set as the final score.
@@ 4.2 Evaluation
Similar to previous work [26], we employ both human evaluation (human-eval) and model-based automatic evaluation (auto-eval). In the former, we use manual grading to judge perfect, acceptable, missing, and incorrect for each answer. In the latter, we merge perfect and acceptable, call it accurate, and use a three-way scoring Scorea with $1,-1,0$ for accurate, incorrect, and missing answers.
We design a two-step method for automatic evaluation: if the answer matches the ground truth exactly, it is considered accurate; otherwise, we use LLMs to determine whether the response is accurate, incorrect, or missing. To avoid the self-preference problem [18], we use two LLM evaluators: ChatGPT (gpt-3.5-turbo[17] and Llama 3 (llama-3-70B-instruct[2] and report the average accurate, hallucination, missing rates, and scores from the two models for each RAG system. Our offline experiment shows that this two-step method yields an average F1 score of $94.7\%$ for ChatGPT and $98.9\%$ for Llama 3 compared to human-eval. See Appendix A.2.2 for more details.
Test data split. We split the data randomly into validation, public test, and private test at 30%, 30%, and 40%, and released the validation and public test sets for the KDD Cup Challenge(Appendix A.2.3).
@@ 5.1 Straightforward RAG solutions
Experiment setup: We started with running LLM-only solutions on the CRAG public test set with $1,335$ questions, using simple prompts that encourage brief answers and “I don’t know” answers when the confidence is low (Appendix A.3.1). We employed Llama 2 Chat (llama-2-7b-chat and llama-2-70b-chat[23], Llama 3 Instruct (llama-3-8B-instruct and llama-3-70B-instruct[2], and GPT-4 Turbo [1]. The web-only RAG solutions we evaluated (Task 1) used a fixed-length web context window (2K tokens for Llama 2 Chat and 4K for Llama 3 Instruct and GPT-4 Turbo); we concatenated webpage snippets using the original order from the data as the reference text, until filling up the window (similar to [25, 11, 16]). Our KG-based solutions (Tasks 2, 3) additionally used a fixed-length KG context window (1K tokens for Llama 2 Chat and 2K for Llama 3 Instruct and GPT-4 Turbo) to include the results from the Mock APIs; we extracted the relevant query entities using llama-3-8B-instruct with in-context learning (similar to [20]) detailed in Appendix A.3.1 and concatenated the results returned from all applicable mock APIs (based on the extracted entities), until filling up the window. We discuss results for Llama 3 70B Instruct and GPT-4 Turbo in this section and give more results for other LLMs in Appendix A.3.2.
Table 5: Performance of straightforward RAG solutions. All numbers are in percentage. LLM-only solutions has up to 34% accuracy and straightforward RAG solutions has up to 44% accuracy.
Model Accuracy Hallucination Missing Scorea
@@
Figure 2: LLM-only and Task 3 solution auto-eval scores (in percentage) across domain, dynamism, popularity, and question type.
@@
Table 5 shows the average evaluation scores from the two auto-evaluators (ChatGPT and Llama 3) and illustrates that the CRAG benchmark is non-trivial. First, the best LLM-only solutions (GPT-4 Turbo) obtained an accuracy of only 34%, with a score of 20%, showing a big room for improvement. Second, straightforward RAG solutions obtained up to 44% accuracy, showing that extra information does help answer more questions reliably. Interestingly, none of the RAG solutions obtain a score higher than 20%; this is because all RAG solutions introduce more hallucinations generated from irrelevant retrieval results, showing a big challenge in RAG—How to judiciously use retrieval results without being distracted by retrieval noises? Third, we found that Task 2 scores are higher than Task 1, showing that the KG knowledge helps improve accuracy, with a similar or even lower hallucination rate, because the KG knowledge is typically brief but precise. Unfortunately, the improvement is mediocre, showing a second challenge in RAG—How to best leverage the power of KG data? Finally, the scores for Task 3 are also higher than Task 2, because of better search ranking (recall that Task 1 and 2 provide five pages randomly selected from the top-10 search results) and better search recall. This shows the importance of search ranking in RAG.
@@
Figure 2 shows the auto-eval scores across the domain, dynamism, popularity, and question type dimension. The results reveal a lot of interesting observations and show that the CRAG benchmark allows more insightful conclusions. First, it shows which slices of the benchmark are harder. For example, we found much lower RAG scores on the Finance and Sports domains, for real-time and fast-changing facts, for tail entities, and for complex questions requiring set answers, post-processing, and with false premises. Second, it shows where it is harder to leverage retrieval results. Taking popularity slices as an example, we observed that GPT-4 Turbo’s score dropped from head (21%) to torso (11%) to tail (8%), consistent with past observations [21]; however, the straightforward RAG solution based on GPT-4 Turbo improved QA quality regarding torso (+7%) and tail entities (+6%) but lowered the quality regarding head (-4%). Finally, although our goal is not to compare different LLMs, the different dimensions allow us to understand the strengths and weaknesses of each method. For example, although the RAG system based on Llama 3 70B Instruct has a lower overall score than the one based on GPT-4 Turbo, it has a similar or slightly higher score in answering simple and comparison questions, whereas much lower scores in answering set and post-processing questions, suggesting investigations on the reasoning capabilities.

### v2

# CRAG – Comprehensive RAG Benchmark

@@ Abstract
Retrieval-Augmented Generation (RAG) has recently emerged as a promising solution to alleviate Large Language Model (LLM)’s deficiency in lack of knowledge. Existing RAG datasets, however, do not adequately represent the diverse and dynamic nature of real-world Question Answering (QA) tasks. To bridge this gap, we introduce the Comprehensive RAG Benchmark (CRAG), a factual question answering benchmark of 4,409 question-answer pairs and mock APIs to simulate web and Knowledge Graph (KG) search. CRAG is designed to encapsulate a diverse array of questions across five domains and eight question categories, reflecting varied entity popularity from popular to long-tail, and temporal dynamisms ranging from years to seconds. Our evaluation of this benchmark highlights the gap to fully trustworthy QA. Whereas most advanced LLMs achieve $\leq 34\%$ accuracy on CRAG, adding RAG in a straightforward manner improves the accuracy only to 44%. State-of-the-art industry RAG solutions only answer $63\%$ of questions without any hallucination. CRAG also reveals much lower accuracy in answering questions regarding facts with higher dynamism, lower popularity, or higher complexity, suggesting future research directions. The CRAG benchmark laid the groundwork for a KDD Cup 2024 challenge and attracted thousands of participants and submissions. We commit to maintaining CRAG to serve research communities in advancing RAG solutions and general QA solutions. CRAG is available at https://github.com/facebookresearch/CRAG/.
@@ 1 Introduction
Large Language Models (LLMs) have transformed the landscape of Natural Language Processing (NLP) tasks, especially in Question Answering (QA) [38, 22, 20, 39]. Despite the advancements, the issue of hallucination persists as a significant challenge; LLMs may generate answers that lack factual accuracy or grounding [27, 14, 30, 32]. Studies have shown that GPT-4’s accuracy in answering questions referring to slow-changing or fast-changing facts is below 15% [36]; even for stable (never-changing) facts, GPT-4’s accuracy in answering questions referring to torso-to-tail (less popular) entities is below 35% [29]. Overcoming hallucinations thus becomes a priority in building reliable QA systems [14, 13].
@@
Retrieval-Augmented Generation (RAG) [12, 19, 6, 8] has recently emerged as a promising solution to alleviate LLM’s deficiency in lack of knowledge and attracted a lot of attention from both academia research and industry. Given a question, a RAG system searches external sources to retrieve relevant information and then provides grounded answers [12, 19, 7] (see Figure 1 for an illustration). Despite its potential, RAG still faces many challenges, such as selecting the most relevant information, reducing question answering latency, and synthesizing information to answer complex questions.
A comprehensive benchmark is currently missing to advance continued research efforts in this field. Our goal is to build a benchmark that can provide a holistic view of the important capabilities and fast but reliable evaluation for RAG to propel the area forward. What is a good benchmark for QA over LLMs? We consider five critical features.
@@
Our second contribution is the evaluation mechanism to allow for reliable comparisons. We designed 3 tasks to test different components in RAG solutions: retrieval summarization, knowledge graph and web retrieval, and end-to-end retrieval-augmented generation (Section 2). Instead of computing the percentage of correctly answered questions, our score system distinguishes hallucinated answers and missing answers, and gives the former a higher penalty as it can be more harmful to ruin user trust. We also design an effective automatic evaluation mechanism to allow for fast evaluations and iterations (Section 4).
Our third contribution is a comprehensive evaluation of straightforward RAG solutions and industry state-of-the-art solutions on RAG (Section 5). Whereas most advanced LLMs achieve $\leq 34\%$ accuracy on CRAG, adding RAG in a straightforward manner improves the accuracy only to 44%. State-of-the-art industry RAG solutions answer only $63\%$ questions without any hallucination, still having much lower accuracy in answering questions regarding facts with higher dynamism, lower popularity, or higher complexity. These evaluations serve two roles: first, they demonstrate that CRAG has appropriate level of difficulty and allows insights drawn from different dimensions of diversities the benchmark has incorporated; second, they highlight the gaps and research directions to a fully trustworthy QA system.
@@
Comparison with existing benchmarks. Table 1 compares CRAG with existing benchmarks for factual question answering. Traditional QA benchmarks such as Natural Questions (NQ) [18], TriviaQA [16], MS MARCO [4], and QALD-10 [35] have advanced QA in the past decade but consider only web retrieved or KG retrieved contents, and do not adequately represent the diverse and dynamic challenges that RAG is facing. New benchmarks for LLM or RAG usually target certain capabilities of the QA system. Researchers created benchmarks to evaluate how well the systems can answer simple knowledge questions [29, 30, 23] and handle more advanced scenarios. These include answering questions with changing answers [36], integrating information from multiple documents [6], addressing multi-hop questions [33], and answering questions with long texts [26]. Moreover, traditional QA benchmarks usually adopt matching-based metrics such as ROUGE [21] or F1 to evaluate the quality of the responses [18, 16]. These metrics, although working well for extractive methods, are known to not perform very effectively for LLMs that generate free-form responses [11].
Despite CRAG being smaller than MS MARCO and NQ, it offers several distinctive advantages: comprehensive coverage, realistic testing with mock APIs, dynamic question handling, diverse fact popularity, extensive content beyond Wikipedia, and fast yet reliable evaluation. These features make CRAG a robust and versatile benchmark for testing RAG systems and broadly QA systems, providing a shared testbed to evaluate how these systems handle real-world, dynamic, and diverse information retrieval and synthesis challenges for reliable LLM-based question answering.
@@
Benchmark Web retrieval KG search Mock API Dynamic question Torso and tail facts Beyond Wikipedia Question size
QALD-10 [35] 0.8K
MS MARCO [4] not explicitly not explicitly 100K
NQ [18] not explicitly not explicitly 323K
RGB [6] 1K
FreshLLM [36] 0.6K
CRAG 4.4K
@@ 2 Problem Description
The three tasks, each adding upon the previous one, allow testing different capabilities of the RAG systems. The only component of a RAG system not covered by these tasks is search retrieval. One may easily extend the tasks to use all 220K webpages in our benchmark as the search corpus for fully end-to-end testing.
@@ 3 Dataset Description
Post-processing-heavy Questions that need reasoning or processing of the retrieved information to obtain the answer (e.g., “how many days did Thurgood Marshall serve as a Supreme Court justice?”).
@@
CRAG contains two parts of data: the QA pairs and the contents for retrieval. We now describe each part of the data. Data generation details can be found in Appendix A.1.1A.1.6.
@@ 3.1 Question answering pairs
Using the above methods, we collected 2,425 Web Questions and 1,984 KG Questions, with $661$, $658$, and $665$ KG Questions containing head, torso, and tail entities respectively. Tables 3 and 4 summarize the distribution of the questions across different dimensions. The size of each dimension slice (e.g., fast-changing facts) allows us to get metrics with $<5\%$ margin-of-error (with 95% confidence level) for most of the cases. The dynamism distribution roughly reflects the nature of the domain (e.g., much more real-time questions for Finance than for other domains). See Appendix A.1.2 for the definition of the dynamism categories.
@@ 3.2 Contents for retrieval
Web search results. For each question, we used the question text as the search query and stored up to $50$ HTML pages from the Brave search API [5]. See Table 8 in Appendix A.1.5 for an example.
We estimated the web search recall with a heuristic-based method: first check whether the ground truth answer URL was found among the pages; if not, decide whether the fact in the ground truths is contained in the page snippet or content with an LLM. In particular, we pass the question, ground truth, and the pages to Llama 3 70B Instruct and ask it to judge if the context is sufficient to answer the question.
Figure 2 shows the estimated web search recall curve for all CRAG questions, with an overall recall of 85% when using all 50 pages. It reflects multiple advantages of the benchmark by design. First, the recall curves are sharp at the beginning and flatten out later, and the recall for the top 5 pages is about 69%. This is comparable to what we observe in practice when developing a RAG system. The non-perfect coverage, especially for Task 1, allows us to test whether the RAG solutions admit “I don’t know” when the retrieval results do not contain the necessary information. Second, compared to the recall from web snippets, full web pages increase the recall by about 20%, emphasizing the importance of extracting and understanding HTML contents. Moreover, the estimated web search recall (50 web pages) is $93\%$ for Web Questions and $74\%$ for KG Questions, indicating significantly lower recall for KG questions than web questions. This aligns with our observations that web search recall for torso and tail entities is typically lower, underlining the crucial role of leveraging KGs in Task 2 and 3.
Figure 2: For $85\%$ of CRAG questions, the web search results are estimated to contain the ground truth facts. The curve shows that the retrieval recall grows sharply at the beginning and flattens out later on.
@@
We collected snapshots of the KG and web search data concurrently while posing real-time and fast-changing questions. This approach ensures that we capture the “snapshot” of the information world at the time of question answering. A RAG solution that performs well on the benchmark should also be capable of reasoning over time and generalizing to evolving questions.
In total, the resulting data contains 220K webpages, a KG of 2.6M entities, and 38 mock APIs. See Table 9 in the Appendix for a complete list of the mock APIs.
@@ 4.1 Metrics
We use a scoring method with score $1$, $0.5$, 00, and $-1$ for each perfect, acceptable, missing, and incorrect answer, respectively, where we penalize hallucinated answers and prefer missing answers to incorrect ones. We then define truthfulness as the average score from all examples in the evaluation set for a given RAG system.
@@ 4.2 Evaluation
Similar to previous work [37], we employ both human evaluation (human-eval) and model-based automatic evaluation (auto-eval). In the former, we use manual grading to judge perfect, acceptable, missing, and incorrect for each answer. In the latter, we merge perfect and acceptable, call it accurate, and use a three-way scoring system with $1,-1,0$ for accurate, incorrect, and missing answers.
We design a two-step method for automatic evaluation: if the answer matches the ground truth exactly, it is considered accurate; otherwise, we use LLMs to determine whether the response is accurate, incorrect, or missing. To avoid the self-preference problem [25], we use two LLM evaluators: ChatGPT (gpt-3.5-turbo-0125[24] and Llama 3 (llama-3-70B-instruct[2] and report the average accurate, hallucination, missing rates, and truthfulness scores from the two models for each RAG system. Our offline experiment shows that this two-step method yields an average F1 score of $94.7\%$ for ChatGPT and $98.9\%$ for Llama 3 compared to human-eval. See Appendix A.2.2 for more details.
Test data split. We split the data randomly into validation (30%), public test (30%), and private (40%), and released the validation and public test sets (Appendix A.2.3). Participants of the KDD Cup challenge can use the validation and public test sets to develop and test their models, and the submitted solutions were evaluated on the private test set. Future offline users of CRAG can use the validation set for development, fine-tuning, and validation, and the public test set for testing and result reporting.
@@ 5.1 Straightforward RAG solutions
Experiment setup: We started with running LLM-only solutions on the CRAG public test set with $1,335$ questions, using simple prompts that encourage brief answers and “I don’t know” answers when the confidence is low (Appendix A.3.1). We employed Llama 2 Chat (llama-2-7b-chat and llama-2-70b-chat[34], Llama 3 Instruct (llama-3-8B-instruct and llama-3-70B-instruct[2], Mixtral (Mixtral-8x7B-Instruct-v0.1[15], Falcon (40B[3], FLAN-T5 (FLAN-T5-XXL[9], and GPT-4 Turbo (gpt-4-turbo-2024-04-09[1]. The web-only RAG solutions we evaluated (Task 1) used a fixed-length web context window (1K tokens for Falcon and FLAN-T5, 2K for Llama 2 Chat, and 4K for Llama 3 Instruct and GPT-4 Turbo); we concatenated webpage snippets using the original order from the data as the reference text, until filling up the window (similar to [36, 17, 23]). Our KG-based solutions (Tasks 2, 3) additionally used a fixed-length KG context window (0.5K tokens for Falcon and FLAN-T5, 1K for Llama 2 Chat, and 2K for Llama 3 Instruct and GPT-4 Turbo) to include the results from the mock APIs; we extracted the relevant query entities using llama-3-8B-instruct with in-context learning (similar to [28]) detailed in Appendix A.3.1 and concatenated the results returned from all applicable mock APIs (based on the extracted entities), until filling up the window. We provide an extensive comparison of all LLMs in Appendix A.3.2 and focus on the best-performing LLMs (i.e., Llama 3 70B Instruct and GPT-4 Turbo) in this section.
Table 5: Performance of straightforward RAG solutions. All numbers are in percentage. LLM-only solutions has up to 34% accuracy and straightforward RAG solutions has up to 44% accuracy. The subscript in “$\text{truthfulness}_{a}$” denotes the result is reported by auto-eval.
Model Accuracy Hallucination Missing Truthfulnessa
@@
Figure 3: LLM-only and Task 3 solution auto-eval truthfulness (in percentage) across domain, dynamism, popularity, and question type.
@@
Table 5 shows the average evaluation results from the two auto-evaluators (ChatGPT and Llama 3) and illustrates that the CRAG benchmark is non-trivial. First, the best LLM-only solutions (GPT-4 Turbo) obtained an accuracy of only 34%, with truthfulness of 20%, showing a big room for improvement. Second, straightforward RAG solutions obtained up to 44% accuracy, showing that extra information does help answer more questions reliably. Interestingly, none of the RAG solutions obtain truthfulness higher than 20%; this is because all RAG solutions introduce more hallucinations generated from irrelevant retrieval results, showing a big challenge in RAG—How to judiciously use retrieval results without being distracted by retrieval noises? Third, we found that Task 2 truthfulness scores are higher than Task 1, showing that the KG knowledge helps improve accuracy, with a similar or even lower hallucination rate, because the KG knowledge is typically brief but precise. Unfortunately, the improvement is mediocre, showing a second challenge in RAG—How to best leverage the power of KG data? Finally, the truthfulness for Task 3 are also higher than Task 2, because of better search ranking (recall that Task 1 and 2 provide five pages randomly selected from the top-10 search results) and better search recall. In particular, we found that the ground truths of over 30% of questions are available in the web retrieval results but are not included in the prompt due to the context window limitation. This shows the importance of search ranking in RAG.
@@
Figure 3 shows the auto-eval results across the domain, dynamism, popularity, and question type dimension. The results reveal a lot of interesting observations and show that the CRAG benchmark allows more insightful conclusions. First, it shows which slices of the benchmark are harder. For example, we found much lower RAG truthfulness on the Finance and Sports domains, for real-time and fast-changing facts, for tail entities, and for complex questions requiring set answers, post-processing, and with false premises. Second, it shows where it is harder to leverage retrieval results. Take the popularity slices as an example, we observed that GPT-4 Turbo’s truthfulness dropped from head (21%) to torso (11%) to tail (8%), consistent with past observations [29]; however, the straightforward RAG solution based on GPT-4 Turbo improved QA quality regarding torso (+7%) and tail entities (+6%) but lowered the quality regarding head (-4%). Finally, although our goal is not to compare different LLMs, the different dimensions allow us to understand the strengths and weaknesses of each method. For example, although the RAG system based on Llama 3 70B Instruct has a lower overall truthfulness score than the one based on GPT-4 Turbo, it has a similar or slightly higher truthfulness in answering simple and comparison questions, whereas much lower truthfulness in answering set and post-processing questions, suggesting investigations on the reasoning capabilities.